OpenStudy (anonymous):

Write the equation of the line that passes through (2,4) and has a slope of -1 in point-slope form. x-4=-1(y-2) x-2=-1(y-4) y-2=-1(x-4) y-4=-1(x-2)

Nnesha (nnesha):

\[\huge\rm y-y_1=m(x-x_1)\] point slope form where m is slope and passes though (x_1,y_1) ponit

Nnesha (nnesha):

point*

imqwerty (imqwerty):

slope intercept form = (y-y1)/(x-x1) = m m=slope (x1,y1) - cordinates of given point so (y-4)/(x-2) = -1 (y-4)=-1(x-2)

Nnesha (nnesha):

that's a formula to find slope slope intercept form is y=mx+b